# Privacy-Enhanced Training-as-a-Service for On-Device Intelligence: Concept, Architectural Scheme, and Open Problems

### v2

@@
Abstract
@@
On-device intelligence (ODI) enables artificial intelligence (AI) applications to run on end devices, providing real-time and customized AI inference without relying on remote servers. However, training models for on-device deployment face significant challenges due to the decentralized and privacy-sensitive nature of users’ data, along with end-side constraints related to network connectivity, computation efficiency, etc. Existing training paradigms, such as cloud-based training, federated learning, and transfer learning, fail to sufficiently address these practical constraints that are prevalent for devices. To overcome these challenges, we propose Privacy-Enhanced Training-as-a-Service (PTaaS), a novel service computing paradigm that provides privacy-friendly, customized AI model training for end devices. PTaaS outsources the core training process to remote and powerful cloud or edge servers, efficiently developing customized on-device models based on uploaded anonymous queries, enhancing data privacy while reducing the computation load on individual devices. We explore the definition, goals, and design principles of PTaaS, alongside emerging technologies that support the PTaaS paradigm. An architectural scheme for PTaaS is also presented, followed by a series of open problems that set the stage for future research directions in the field of PTaaS.
Index Terms:
On-device intelligence, service computing, privacy computing
I Introduction
@@
On-device intelligence (ODI) [1] is an emerging technology that combines mobile computing and artificial intelligence (AI), empowering end-side on-device models to deliver real-time, customized intelligent inference without network connectivity. ODI shows considerable potential in the forthcoming Internet of Everything (IoE) era, as evidenced by applications such as intelligent medical diagnosis, AI-enhanced motion tracking [2], and behavior fingerprinting [3]. Leading technology corporations such as Google, Anthropic, and Xiaomi have recognized ODI as an essential technology in their flagship products, represented by Gemini-Nano [4], Claude-3-Haiku [5], and MiLM-1.3B [6].
Figure 1: Schematic diagram of PTaaS.
@@
Despite these advancements, ODI faces significant challenges due to the decentralized and privacy-sensitive nature of user data derived from the network termination. Hence, it is difficult for a single device to train a powerful AI model independently, especially with the massive end-side limitations related to network connectivity, computation capability, etc [2]. To navigate these constraints and harness the full potential of ODI, various methods have been proposed to seek a balance between the need for powerful, customized AI training services and the practical constraints inherent in end devices.
@@
In the following sections, we provide a comprehensive elaboration on PTaaS from the perspectives of definition, goals, and design principles. Besides, we analyze the emerging technologies that support PTaaS, including privacy computing, cloud-edge collaboration, transfer learning, and information retrieval. Furthermore, we provide the PTaaS hierarchy structure consisting of five layers: infrastructure, data, algorithms, services, and applications. Finally, we discuss a series of open problems faced by PTaaS in terms of improving privacy protection mechanisms, cloud-edge resource collaborative management, optimization of customized model training, designing pricing strategies, and standard specifications.
II Concept of Privacy-Enhanced Training-as-a-service
II-A Definition of PTaaS
@@ II-B Goals of PTaaS
II-C Design Principles of PTaaS
@@
Figure 2: Overview of the five-layer hierarchy structure of PTaaS.
III Architectural Scheme of Privacy-Enhanced Training-as-a-service
III-A Hierarchy Structure of PTaaS
@@
III-B Emerging Technologies Related to PTaaS
@@
Figure 3: Framework of the instantiated PTaaS process based on cloud computing.
III-C PTaaS Process Instantiation
To illustrate the on-demand training services offered by PTaaS, we provide a streamlined explanation of a simplified PTaaS process based on cloud computing, as shown in Fig. 3. This process demonstrates how PTaaS can be applied to image classification for smartphones.
@@ III-C PTaaS Process Instantiation
  • Customized queries delivery. The smartphones further encrypt the anonymous hash vectors, metadata, as well as users’ customized image classification model (collectively referred to as queries), and then upload the queries to the cloud. The cloud decrypts and verifies the uploaded queries, refining users’ model training requirements to ensure optimal trained results.
@@
  • Model offline deployment. The trained models are transmitted back to users’ smartphones, where they undergo an integrity check before being deployed to the local inference engine for practical inference tasks. This offline deployment ensures that the model can operate efficiently on devices without requiring cloud connectivity.
@@
IV Open Problems of Privacy-Enhanced Training-as-a-service
@@
V Conclusions and Future Works
@@
Acknowledgement
We thank GPT-4 and Claude-3 for their invaluable contributions as brainstorming partners and their assistance in language refinement. We also thank professor Zhuotao Liu from Tsinghua University for inspiring suggestions.

### v3

@@
Privacy-Enhanced Training-as-a-Service for On-Device Intelligence: Concept, Architectural Scheme, and Open Problems
Zhiyuan Wu1,2, Sheng Sun1, Yuwei Wang1, Min Liu1,3, Bo Gao4, Tianliu He1,2, Wen Wang1,2
{CJK*}
GBKsong
1Institute of Computing Technology, Chinese Academy of Sciences
2University of Chinese Academy of Sciences
3Zhongguancun Laboratory
4Beijing Jiaotong University
Contact Email: ywwang@ict.ac.cn
footnotetext:
 Regular Paper
@@
AbstractOn-device intelligence (ODI) enables artificial intelligence (AI) applications to run on end devices, providing real-time and customized AI inference without relying on remote servers. However, training models for on-device deployment face significant challenges due to the decentralized and privacy-sensitive nature of users’ data, along with end-side constraints related to network connectivity, computation efficiency, etc. Existing training paradigms, such as cloud-based training, federated learning, and transfer learning, fail to sufficiently address these practical constraints that are prevalent for devices. To overcome these challenges, we propose Privacy-Enhanced Training-as-a-Service (PTaaS), a novel service computing paradigm that provides privacy-friendly, customized AI model training for end devices. PTaaS outsources the core training process to remote and powerful cloud or edge servers, efficiently developing customized on-device models based on uploaded anonymous queries, enhancing data privacy while reducing the computation load on individual devices. We explore the threat model, definition, goals, and design principles of PTaaS, alongside emerging technologies that support the PTaaS paradigm. An architectural scheme for PTaaS is also presented, followed by a series of open problems that set the stage for future research directions in the field of PTaaS.
KeywordsOn-device intelligence, Service computing, Privacy computing
1 Introduction
@@
On-device intelligence (ODI) [1, 2] is an emerging technology that combines mobile computing and artificial intelligence (AI), empowering end-side on-device models to deliver real-time, customized intelligent inference without network connectivity. ODI shows considerable potential in the forthcoming Internet of Everything (IoE) era, as evidenced by applications such as intelligent medical diagnosis, AI-enhanced motion tracking [3], and behavior fingerprinting [4]. Leading technology corporations such as Google, Anthropic, and Xiaomi have recognized ODI as an essential technology in their flagship products, represented by Gemini-Nano [5], Claude-3-Haiku [6], and MiLM-1.3B [7].
@@
Despite these advancements, ODI faces significant challenges due to the decentralized and privacy-sensitive nature of user data derived from the network termination [8, 9]. Hence, it is difficult for a single device to train a powerful AI model independently, especially with the massive end-side limitations related to network connectivity, computation capability, etc [3]. To navigate these constraints and harness the full potential of ODI, various methods have been proposed to seek a balance between the need for powerful, customized AI training services and the practical constraints inherent in end devices.
Fig. 1: Schematic diagram of PTaaS.
@@
In the following sections, we provide a comprehensive elaboration on PTaaS from the perspectives of the threat model, definition, goals, and design principles. Besides, we analyze the emerging technologies that support PTaaS, including privacy computing, cloud-edge collaboration, transfer learning, and information retrieval. Furthermore, we provide the PTaaS hierarchy structure consisting of five layers: infrastructure, data, algorithms, services, and applications. Finally, we discuss a series of open problems faced by PTaaS in terms of improving privacy protection mechanisms, cloud-edge resource collaborative management, optimization of customized model training, designing pricing strategies, and standard specifications.
2 Threat Model
We establish a threat model that characterizes the security assumptions and potential vulnerabilities within the PTaaS service paradigm. The model encompasses the following key entities and their security properties:
  • Remote servers. The remote servers (both cloud and edge) are assumed to be honest but curious [17], meaning they faithfully execute the prescribed protocols but may attempt to infer sensitive information from the received data. While the servers are authorized to decrypt uploaded information for computation purposes, they are not trusted to access raw data generated on end devices.
  • Communication channels. The communication channels between devices and servers are vulnerable to malicious third-party attackers who may attempt to eavesdrop on or tamper with transmitted data [18]. These adversaries are assumed to have the capability to monitor and intercept all information passing through the network infrastructure.
  • End devices. End devices are trusted to generate and maintain data locally while executing assigned tasks. However, sharing raw local data on devices could cause inadvertent leakage of sensitive information, which should be strictly prohibited [9].
This threat model guides the design of PTaaS security mechanisms, ensuring robust privacy protection while maintaining system functionality.
3 Concept Elaboration
3.1 Definition of PTaaS
@@ 3.2 Goals of PTaaS
3.3 Design Principles of PTaaS
@@
Fig. 2: Overview of the five-layer hierarchy structure of PTaaS.
4 Architectural Scheme
4.1 Hierarchy Structure of PTaaS
@@
4.2 Emerging Technologies Related to PTaaS
@@
Fig. 3: Framework of the instantiated PTaaS process based on cloud computing.
Algorithm 1 Privacy-Enhanced Training-as-a-Service (PTaaS).
procedure DeviceExecute()
      //iterative model updates
       while require_model_update is true:
             //data collection and anonymization
             device_data = collect_data(end_device)
             device_model = create_model(end_device)
             anonymized_data = hash_function(device_data)
             differentially_private_anonymized_data = add_differential_privacy_noise(anonymized_data)
             //customized queries delivery
             query = encrypt(differentially_private_anonymized_data, device_model)
             transmit query to the cloud
             call CloudExecute()
             //model offline deployment
             receive encrypted_trained_device_model from the cloud
             if check_integrity(encrypted_trained_device_model) is true:
                   trained_device_model = decrypt(encrypted_trained_device_model)
                   deploy_to_inference_engine(trained_device_model)
end procedure
procedure CloudExecute()
       //customized queries delivery
       receive query from end device
       if verify_query(query) is true:
             differentially_private_anonymized_data, device_model = decrypt(encrypted_query)
             //cloud_data_retrieval
             relevant_data = similarity_search(differentially_private_anonymized_data, cloud_dataset)
             //customized model training
             trained_device_model = transfer_learning(device_model, relevant_data, pretrained_model)
             //model offline deployment
             encrypted_trained_device_model = encrypt(trained_device_model)
             transmit encrypted_trained_device_model to device
end procedure
4.3 PTaaS Process Instantiation
To illustrate the on-demand training services offered by PTaaS, we provide a streamlined explanation of a simplified PTaaS process based on cloud computing, as shown in Fig. 3 and Algorithm 1. This process demonstrates how PTaaS can be applied to image classification for smartphones.
@@ 4.3 PTaaS Process Instantiation
  • Customized queries delivery. The smartphones further encrypt the anonymous hash vectors, metadata, as well as users’ customized image classification model (collectively referred to as queries), and then upload the queries to the cloud. The cloud verifies and decrypts the uploaded queries, refining users’ model training requirements to ensure optimal trained results.
@@
  • Model offline deployment. The trained models are encrypted and transmitted back to users’ smartphones, where they undergo decryption and an integrity check before being deployed to the local inference engine for practical inference tasks. This offline deployment ensures that the model can operate efficiently on devices without requiring cloud connectivity.
@@
5 Open Problems
@@
6 Discussion
6.1 Security Mechanisms in PTaaS
The security of PTaaS relies on three fundamental mechanisms: sensitive hashing, differential privacy, and encryption schemes, which work together to protect private on-device data during the service process. For sensitive hashing, MinHash [21] and SimHash [22] can be applied for data anonymization and similarity-based retrieval, with MinHash ideal for set-based data, and SimHash excels in processing high-dimensional local data. For differential privacy, the Laplace Mechanism [23] is a standard choice for numerical anonymous data. In scenarios where the anonymous data follows a Gaussian distribution, the Gaussian Mechanism [24] is better suited due to its compatibility with the inherent statistical properties. For encryption, AES [25] provides a robust balance between security guarantees and computational efficiency, making it well-suited for resource-constrained devices.
6.2 Feasibility Analysis of PTaaS
The feasibility of PTaaS can be evaluated from two key perspectives: technological support and practical applications. From a technological perspective, PTaaS leverages several well-established technologies, including cloud computing, privacy computing, and transfer learning, which collectively provide a robust foundation for its implementation. The proposed five-layer hierarchical architecture further facilitates the modular development and scalability of PTaaS platforms, ensuring flexibility and adaptability to evolving technological trends. From a practical perspective, PTaaS addresses critical real-world challenges in ODI deployment. Its ability to deliver privacy-enhanced, customized model training while offloading computational burdens from resource-constrained devices makes it particularly suitable for mobile scenarios with limited computational power, energy, or storage. This aligns with the growing demand for on-device AI applications across diverse environments. Moreover, the simplified one-shot query mechanism significantly lowers the barrier to entry for users, offering a more accessible alternative to existing approaches, such as federated learning and traditional transfer learning, which often require extensive communication or local computation. These advantages make PTaaS a promising paradigm for the widespread adoption of ODI-enabled applications.
6.3 Privacy Level of PTaaS
While PTaaS enhances device-side privacy in outsourcing the core training process, it is important to clarify the level of privacy protection it provides. By requiring devices to share only anonymous information derived from local data, PTaaS avoids the transmission of raw user data over the network, thereby delivering a meaningful improvement in privacy protection. Despite these superiorities, PTaaS enhances privacy rather than preserves it completely. Since service providers must decrypt and process certain anonymized information to train models, there remains a theoretical risk of privacy breaches. For instance, adversaries could potentially reconstruct private information by optimizing for hash similarity. As such, while PTaaS significantly improves data privacy compared to centralized training paradigms, it is not impervious to all privacy risks.
7 Conclusions and Future Works